# Synthetic bootstrapped pretraining

### v2

@@ Abstract
We introduce Synthetic Bootstrapped Pretraining (SBP), a language model (LM) pretraining procedure that first learns a model of relations between documents from the pretraining dataset and then leverages it to synthesize a vast new corpus for joint training. While the standard pretraining teaches LMs to learn causal correlations among tokens within a single document, it is not designed to efficiently model the rich, learnable inter-document correlations that can potentially lead to better performance. We validate SBP by designing a compute-matched pretraining setup and pretrain a 3B-parameter model on up to 1T tokens from scratch. We find SBP consistently improves upon a strong repetition baseline and delivers a significant fraction of performance improvement attainable by an oracle upper bound with access to 20x more unique data. Qualitative analysis reveals that the synthesized documents go beyond mere paraphrases – SBP first abstracts a core concept from the seed material and then crafts a new narration on top of it. Besides strong empirical performance, SBP admits a natural Bayesian interpretation: the synthesizer implicitly learns to abstract the latent concepts shared between related documents.
@@ 1 Introduction
Pretraining on the diverse internet texts is now seen to be bottlenecked by the rapid depletion of high-quality text data (Villalobos et al., 2024). This imminent “scaling wall” motivates us to utilize existing data more effectively. Re-examining the conceptual foundation of pretraining, its success originates from the rich causal correlation among tokens within a document. However, this is not the only source of correlation pretraining dataset contains: a code document implementing the attention mechanism is derived from the arXiv preprint of the transformer paper; The book of Harry Potter is structurally similar to the screenplay of its movie production. Such connections suggest a weaker form of inter-document correlation derived from an underlying joint distribution of pretraining documents. We hypothesize that this additional signal, which is missed by the standard pretraining, can be captured by synthetic data, presenting an underexplored avenue for improving performance.
@@
To test our hypothesis, we design a compute-matched, data-constrained experimental framework under which we pretrain a 3B-parameter model on up to 1T tokens from scratch (Li et al., 2024; Zyphra, 2024), demonstrating the potential applicability of SBP for advancing frontier LMs. We compare SBP’s performance against two crucial references: a strong repetition baseline, which represents the standard approach in data-constrained settings, and an oracle upper bound, which has access to an unlimited pool of unique internet data (§3). Our results show that SBP consistently surpasses the strong repetition baseline across different pretraining scales and closes a significant portion of the performance gap to the oracle with 20x additional unique data access (§4.1).
@@
  • Large-scale empirical validation: We design a compute-matched pretraining setup that enables rigorous measurement of LM self-improvement and empirically validate SBP on a 3B-parameter model trained on up to 1T tokens from scratch.
@@ 1.1 Related work
The concept of pretraining, closest to its modern form, originates from a series of works including ELMo (Peters et al., 2018), ULMFiT (Howard & Ruder, 2018), BERT (Devlin et al., 2019), that propose to pretrain a neural network via an unsupervised objective and subsequently finetune for a wide range of downstream tasks. The GPT-series (Radford et al., 2018a; b; Brown et al., 2020; OpenAI et al., 2024) cemented the practice of using next-token prediction as the pretraining objective and applying it to large-scale crawled webpages as opposed to task-specific datasets (e.g., English-to-French translation). In recent years, the size of the pretraining corpora has grown rapidly, driven by the availability of massive web-crawled datasets, leading to a successful stream of dataset and pretrained model artifact: BERT (Devlin et al., 2019; Liu et al., 2020), GPT-2 WebText (Radford et al., 2018b), CommonCrawl (Common Crawl, 2007), CCNet (Wenzek et al., 2019), T5 C4 (Raffel et al., 2020), the Pile (Gao et al., 2020), Gopher Massive Text (Rae et al., 2021), Llamda series (Touvron et al., 2023; Dubey et al., 2024), RefinedWeb (Penedo et al., 2023), Dolma (Soldaini et al., 2024), DCLM-baseline (Li et al., 2024), NemotronCC (Su et al., 2024), etc. While pretraining has been tremendously successful, the rapid depletion of available internet text motivates us to shift our focus from acquiring more data to using the existing data more effectively.
@@
A natural way to overcome the limitations of scarce high-quality web data is to pretrain (Gunasekar et al., 2023; Abdin et al., 2023; 2024a; 2024b; Team et al., 2025) or continually pretrain (Yang et al., 2025; Ruan et al., 2025; Zweiger et al., 2025; Nguyen et al., 2025) LMs on synthetic data. Existing approaches to data synthesis rely on distillation from a powerful “teacher” LM that generates compressed knowledge representation for the “student” LM to learn (Hinton et al., 2015). These teacher models must first undergo a human alignment process, which requires extensive human annotations and preference data (Ouyang et al., 2022). Synthetic data from the teacher LM hints at a limited scaling trend: whilst the synthesized data from the teacher LM can be as impressive (DatologyAI et al., 2025) as 7x more effective than real data, the performance improvement quickly converges to that of the teacher LM (Busbridge et al., 2025). We instead consider the scenario where the sole source of world knowledge comes from a fixed set of pretraining documents (e.g., the internet) and algorithmically learn a data synthesizer with minimal human intervention (e.g., generative teacher models or human writing prompts). Therefore, our experiment setup simulates a situation where the LMs can self-boost their pretraining capability by refining their understanding of the fixed collection of pretraining documents.
@@
A natural class of methods that incorporates multiple documents together is retrieval augmented generation (RAG) (Lample et al., 2019; Lewis et al., 2020). While originally introduced as a technique to be used at test-time for a domain-specific downstream task (Borgeaud et al., 2021; Li et al., 2022), retrieval augmented approaches have been extended in scope: Khandelwal et al. (2020) and Yang et al. (2023) implement RAG at pretraining scale and show improved test perplexity; Guu et al. (2020) incorporates RAG at pretraining time by jointly training a retriever and the model itself for improved QA performance. Shi et al. (2024) groups related documents into the same context window for improved long-context capability. In general, while the RAG-related approach enables the model to utilize rich inter-document correlations, it is fundamentally limited by the context window of the LM. In contrast, SBP encodes correlations into synthetic data that can be iteratively learned by the LM one document at a time. Prior to the advancement of embedding models that allow for retrieving the entire document, Guu et al. (2018) proposed retrieving neighboring pairs of sentences using Jaccard similarity and modeling the conditional distribution between them, which is similar to our conditional data synthesizer objective; however, they did not perform any pretraining experiments.
@@ 2.1 Data-contrained pretraining setup
We consider a data-constrained setup where the goal is to train the best-performing LM given access to a fixed document collection ${\mathcal{D}}_{\text{pretrain}}$} (e.g., a snapshot of the entire internet). To establish a controlled experimental framework, we also choose a transformer architecture with parameters $\theta$ and a collection of held-out evaluation benchmarks $\mathrm{Perf}$} (e.g., perplexity, few-shot QA accuracy). Recall that a transformer takes in a sequence of tokens and outputs a sequence of conditional probabilities of each token given all previous tokens. Applying the chain rule for joint probability, we can use a transformer to calculate the probability $p_{\theta}(x)$ of observing a particular text input $x$, or the conditional probability $p_{\theta}(x|y)$ of one text $x$ given another $y$.
@@
$\operatorname*{arg\,max}_{\theta}\sum_{{d}\in{{\mathcal{D}}_{\text{pretrain}}}}\log p_{\theta}({d}).$ (1)
@@ 2.2 Synthetic bootstrapped pretraining
In preparation for training the conditional data synthesizer, SBP first curates pairs of related documents. To efficiently perform similarity search at pretraining scale, we adopt the Approximate Nearest Neighbor (ANN) methodology (Malkov & Yashunin, 2018), which embeds each document as a quantized vector normalized to the unit sphere and then performs massively parallelizable linear algebraic operations. In our implementation of SBP, we use inner-product similarity, which we denote by $\langle{d_{1}},{d_{2}}\rangle$. Then, we select a subset of pairs whose similarity score exceeds a certain threshold $\alpha$:
@@ 3 Experiment setup
In this section, we present our concrete experimental implementation of SBP. In a nutshell, we curated a pretraining dataset of 582M high-quality documents totaling 482B tokens from DCLM (Li et al., 2024), designed a 3B-parameter transformer architecture modified from the Llama 3 implementation (Dubey et al., 2024), and chose nine commonly used benchmarks targeted at general world knowledge and commonsense reasoning (§3.1). We propose a compute-matched comparison scheme to validate SBP against natural reference methods at a compute scale of up to 1T total training tokens in our largest experiment (§3.2), bringing validation at a scale relevant for frontier LM development.
@@ 3.1 Data, model, and evaluation
A typical pretraining dataset is a mixture of different sources (e.g., GitHub, arXiv, CommonCrawl, etc.) with distinct sampling weights assigned to each constituent. We simplify this reality by considering a fixed document collection, which is a customized version of the DCLM dataset (Li et al., 2024). The original 4T token DCLM-baseline split contains roughly 80% duplicates, as reported by Zyphra (2024). Therefore, we begin with the de-duplicated dataset, which consists of 769B tokens. We clean the raw Zyphra de-duplicated data by normalizing repeated line breaks, removing long URL links, and fixing malformed Unicode characters. For efficiency reasons, we cap the context window of the synthesizer-tuning (3) step at 8,192 tokens. As a result, we additionally filter out the documents whose length is above 4,096 tokens, allowing both $d_{1}$} and $d_{2}$} to fit into the context window in the worst case when both documents are 4,096 tokens long. After all the de-duplication, cleaning, and filtering procedures, we end up with a collection of 582M high-quality documents ${\mathcal{D}}_{\text{pretrain}}$} totaling 482B tokens. We use the notation $|{{\mathcal{D}}_{\text{pretrain}}}|$ to denote the number of documents in the pretraining dataset and $\|{{\mathcal{D}}_{\text{pretrain}}}\|$ to denote the total number of tokens.
@@
We use the Llama 3 transformer architecture (Dubey et al., 2024) to model the probability $p_{\theta}$ with the notable exception of implementing a QK-norm on top of the existing design, which we empirically find to stabilize training. Our resulting model is a 3B-parameter 26-layer transformer model with a hidden dimension of 3,072. Each layer employs grouped query attention with 24 query heads and 8 key/value heads. The position embedding is RoPE (Su et al., 2023) for queries and keys, with frequency 5e+5. The feedforward network (FFN) has hidden dimension 8,064, and we apply prenorm to both the attention and FFN blocks. For tokenization, we implement a customized BPE tokenization with a vocabulary size of 49,152. To match the 8,192 context window design for synthesizer-tuning we have mentioned, we use context window 4,096 for pretraining, so that every document in ${\mathcal{D}}_{\text{pretrain}}$} can fit into the context window.
@@
To assess the pretraining capability of LM, we measure pretraining test loss and general world knowledge benchmarks. We evaluate held-out test perplexity (exponential of negative log-probability) on 1) OpenWebText2 from EleutherAI (Radford et al., 2018b); 2) Narrative understanding with LAMBADA (Paperno et al., 2016) and 3) Broad domain multiple-choice with MMLU (Hendrycks et al., 2020). We evaluate QA accuracy on 4) Hard scientific reasoning with ARC‑Challenge (Clark et al., 2018); 5) Easy scientific reasoning with ARC‑Easy (Clark et al., 2018); 6) Scientific QA with SciQ (Welbl et al., 2017); 7) Common sense reasoning with Winogrande (Sakaguchi et al., 2021); 8) Reading comprehension with TriviaQA (Joshi et al., 2017); 9) Openbook QA with WebQS (Berant et al., 2013). We directly evaluate the pretrained model with either zero-shot or few-shot prompts. Although MMLU is more commonly known as a QA benchmark, we find that evaluating MMLU accuracy for weak models yields a highly non-smooth readout. As a result, for each MMLU test question, we prepend the question with a 5-shot example of QA pairs and postpend it with the correct answer. Then, we treat each such sample as a text corpus and evaluate LM’s perplexity on such a text sample. Empirically, we find that this perplexity-based MMLU correlates well with MMLU accuracy when the underlying model is large enough to yield a stable readout, and also delivers smooth performance changes for smaller models. Note that those benchmarks are known to improve significantly with instruction finetuning (Wei et al., 2022). However, we stick to our data-constrained setup and do not introduce any additional data that may confound the comparison.
@@ 3.2 Compute-matched comparsion
We propose a compute-matched experimentation framework to rigorously compare SBP against two natural references: a repetition baseline where we repeat ${\mathcal{D}}_{\text{pretrain}}$} multiple times to utilize the available training compute and an oracle upper bound that enables the model to access as many unique documents as possible. Operationally, we control the training compute by controlling the total tokens seen during training, which is proportional to the training FLOPs given a fixed batch size and context window. We validate SBP across two different scales:
@@
  • 1T-scale: We also consider a larger scale closer to frontier model training, where we cap the training compute at 1T tokens and data access at $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B tokens.
@@
Since the compute budget typically exceeds the total number of unique tokens $\|{{\mathcal{D}}_{\text{pretrain}}}\|$, a natural baseline to use the additional compute is to repeat ${\mathcal{D}}_{\text{pretrain}}$} over multiple epochs. By design, in both 200B-scale and 1T-scale, we repeat the pretraining dataset ${\mathcal{D}}_{\text{pretrain}}$} 20 times to exploit the available compute budget. In practice, when the pretraining dataset comes from a mixture of different sources, higher-quality documents can be seen as many as 30 times during pretraining, while lower-quality texts may appear only once. Muennighoff et al. (2023) systematically evaluates the repetition baseline as a proposal to scale LMs under data constraints and finds that repeating ${\mathcal{D}}_{\text{pretrain}}$} up to 4 times yields nearly no performance degradation compared with having access to unlimited fresh data, but after around 40 times, repetition yields rapidly diminishing returns. Therefore, our choice of 20 times repetition with compute-matched comparison strikes a reasonable balance between efficient experimental execution and exhausting all possible performance gains from a fixed ${\mathcal{D}}_{\text{pretrain}}$} via repetition.
@@
Besides showing improvement against the repetition baseline, we also evaluate an oracle upper bound with unlimited data access. The motivation behind this is to contextualize the numerical improvement delivered by SBP. As we shall see in the next section, because different benchmarks respond differently to data size changes, SBP can deliver an improvement as large as 3.74% on some benchmarks but only 0.14% on others (Table 1). Also, as performance on LM benchmarks tend to scale logarithmically (Owen, 2024; Kaplan et al., 2020) against data improvement, the numerical difference quickly caps out as we move from the 200B scale to the 1T-scale. By introducing this oracle upper bound, we can contrast the SBP improvement against this “oracle” improvement.
@@
For both the repetition baseline and oracle upper bound at both 200B-scale and 1T-scale, we use a batch size of 2,048 and a context window of 4,096, resulting in a throughput of 8M tokens per step. We apply a cosine learning rate scale with a 5% warmup to a peak learning rate of 1e-2, followed by subsequent decay to 5e-5 towards the end. Under this setup, pretraining costs 11K v5p-TPU hours at 200B-scale and 59K v5p-TPU hours at 1T-scale. For a clean comparison, we adhere to this hyperparameter throughout the paper, including the SBP experiment presented next.

### v3

@@ Abstract
We introduce Synthetic Bootstrapped Pretraining (SBP), a language model (LM) pretraining procedure that first learns a model of relations between documents from the pretraining dataset and then leverages it to synthesize a vast new corpus for joint training. While the standard pretraining teaches LMs to learn causal correlations among tokens within a single document, it is not designed to efficiently model the rich, learnable inter-document correlations that can potentially lead to better performance. We validate SBP by designing a compute-matched pretraining setup and pretrain a 3B-parameter and a 6B-parameter model on up to 1T tokens from scratch. We find SBP consistently improves upon a strong repetition baseline and delivers up to 60% of performance improvement attainable by an oracle upper bound with access to 20x more unique data. Qualitative analysis reveals that the synthesized documents go beyond mere paraphrases – SBP first abstracts a core concept from the seed material and then crafts a new narration on top of it. Besides strong empirical performance, SBP admits a natural Bayesian interpretation: the synthesizer implicitly learns to abstract the latent concepts shared between related documents.
@@ 1 Introduction
Pretraining on the diverse internet texts is now seen to be bottlenecked by the rapid depletion of high-quality text data (villalobos2024run). This imminent “scaling wall” motivates us to utilize existing data more effectively. Re-examining the conceptual foundation of pretraining, its success originates from the rich causal correlation among tokens within a document. However, this is not the only source of correlation pretraining datasets contain: a code document implementing the attention mechanism is derived from the arXiv preprint of the transformer paper; The book of Harry Potter is structurally similar to the screenplay of its movie production. Such connections suggest a weaker form of inter-document correlation derived from an underlying joint distribution of pretraining documents. We hypothesize that this additional signal, which is missed by standard pretraining, can be captured by synthetic data, presenting an underexplored avenue for improving performance.
@@
To test our hypothesis, we design a compute-matched, data-constrained experimental framework under which we pretrain a 3B-parameter and a 6B-parameter model on up to 1T tokens from scratch (li2024datacomplm; zyphradedup), demonstrating the potential applicability of SBP for advancing frontier LMs. We compare SBP’s performance against two crucial references: a strong repetition baseline, which represents the standard approach in data-constrained settings, and an oracle upper bound, which has access to an unlimited pool of unique internet data (§3). Our results show that SBP consistently surpasses the strong repetition baseline across different pretraining scales and closes up to 60% of the performance gap to the oracle with 20x additional unique data access (§4.1).
@@
  • Large-scale empirical validation: We design a compute-matched pretraining setup that enables rigorous measurement of LM self-improvement and empirically validate SBP on 3B and 6B parameter models trained on up to 1T tokens from scratch.
@@ 1.1 Related work
The concept of pretraining, closest to its modern form, originates from a series of works including ELMo (peters2018deepcontextualizedwordrepresentations), ULMFiT (howard2018universallanguagemodelfinetuning), BERT (devlin-etal-2019-bert), that propose to pretrain a neural network via an unsupervised objective and subsequently finetune for a wide range of downstream tasks. The GPT-series (gpt1; gpt2; gpt3; gpt4) cemented the practice of using next-token prediction as the pretraining objective and applying it to large-scale crawled webpages as opposed to task-specific datasets (e.g., English-to-French translation). In recent years, the size of the pretraining corpora has grown rapidly, driven by the availability of massive web-crawled datasets, leading to a successful stream of dataset and pretrained model artifact: BERT (devlin-etal-2019-bert; liu2020roberta), GPT-2 WebText (gpt2), CommonCrawl (commoncrawl), CCNet (wenzek2019ccnet), T5 C4 (t5), the Pile (gao2020pile), Gopher Massive Text (rae2021scaling), Llamda series (touvron2023llama2openfoundation; llama3), RefinedWeb (penedo2023refinedweb), Dolma (soldaini2024dolma), DCLM-baseline (li2024datacomplm), NemotronCC (su2024nemotron), etc. While pretraining has been tremendously successful, the rapid depletion of available internet text motivates us to shift our focus from acquiring more data to using the existing data more effectively.
@@
A natural way to overcome the limitations of scarce high-quality web data is to pretrain (phi1; phi2; phi3; phi4; kimiteam2025kimik2openagentic) or continually pretrain (entigraph; ruan2025reasoning; zweiger2025selfadaptinglanguagemodels; nguyen2025recyclingwebmethodenhance) LMs on synthetic data. Existing approaches to data synthesis rely on distillation from a powerful “teacher” LM that generates compressed knowledge representation for the “student” LM to learn (hinton2015distillingknowledgeneuralnetwork). These teacher models must first undergo a human alignment process, which requires extensive human annotations and preference data (instruct_gpt). Synthetic data from the teacher LM hints at a limited scaling trend: whilst the synthesized data from the teacher LM can be as impressive (datologyai2025beyondweblessonsscalingsynthetic) as 7x more effective than real data, the performance improvement quickly converges to that of the teacher LM (busbridge2025distillation). We instead consider the scenario where the sole source of world knowledge comes from a fixed set of pretraining documents (e.g., the internet) and algorithmically learn a data synthesizer with minimal human intervention (e.g., generative teacher models or human writing prompts). Therefore, our experiment setup simulates a situation where the LMs can self-boost their pretraining capability by refining their understanding of the fixed collection of pretraining documents.
@@
A natural class of methods that incorporates multiple documents together is retrieval augmented generation (RAG) (Lample:2019; rag). While originally introduced as a technique to be used at test-time for a domain-specific downstream task (Borgeaud:2021; li2022decoupled), retrieval augmented approaches have been extended in scope: Khandelwal:2020 and resmem implement RAG at pretraining scale and show improved test perplexity; Guu:2020 incorporates RAG at pretraining time by jointly training a retriever and the model itself for improved QA performance. shi2024incontext groups related documents into the same context window for improved long-context capability. In general, while the RAG-related approach enables the model to utilize rich inter-document correlations, it is fundamentally limited by the context window of the LM. In contrast, SBP encodes correlations into synthetic data that can be iteratively learned by the LM one document at a time. Prior to the advancement of embedding models that allow for retrieving the entire document, guu2018generatingsentenceseditingprototypes proposed retrieving neighboring pairs of sentences using Jaccard similarity and modeling the conditional distribution between them, which is similar to our conditional data synthesizer objective; however, they did not perform any pretraining experiments.
@@ 2.1 Data-contrained pretraining setup
We consider a data-constrained setup where the goal is to train the best-performing LM given access to a fixed document collection ${\mathcal{D}}_{\text{pretrain}}$} (e.g., a snapshot of the entire internet). To establish a controlled experimental framework, we also choose a transformer architecture with parameters $\theta$ and a collection of held-out evaluation benchmarks $\mathrm{Perf}$} (e.g., perplexity, few-shot QA accuracy). Recall that a transformer takes in a sequence of tokens and outputs a sequence of conditional probabilities of each token given all previous tokens. Applying the chain rule for joint probability, we can use a transformer to calculate the probability $p_{\theta}(y)$ of observing a particular text input $y$, or the conditional probability $p_{\theta}(y|x)$ of one piece of text $y$ followed by $x$.
@@
$\operatorname*{arg\,max}_{\theta}\sum_{{d}\in{{\mathcal{D}}_{\text{pretrain}}}}\log p_{\theta}({d}),$ (1)
@@ 2.2 Synthetic bootstrapped pretraining
In preparation for training the conditional data synthesizer, SBP first curates pairs of related documents. To efficiently perform similarity search at pretraining scale, we adopt the Approximate Nearest Neighbor (ANN) methodology (malkov2018efficientrobustapproximatenearest), which embeds each document as a quantized vector normalized to the unit sphere and then performs massively parallelizable linear algebraic operations. In our implementation of SBP, we use inner-product similarity, which we denote by $\langle{d_{1}},{d_{2}}\rangle$. Then, we select a subset of pairs whose similarity score exceeds a certain threshold $\alpha$:
@@ 3 Experiment setup
In this section, we present our concrete experimental implementation of SBP. In a nutshell, we curated a pretraining dataset of 582M high-quality documents totaling 482B tokens from DCLM (li2024datacomplm), designed 3B and 6B transformer architectures modified from the Llama 3 implementation (llama3), and chose nine commonly used benchmarks targeted at general world knowledge and commonsense reasoning (§3.1). We propose a compute-matched comparison scheme to validate SBP against natural reference methods at a compute scale of up to 1T total training tokens in our largest experiment (§3.2), bringing validation at a scale relevant for frontier LM development.
@@ 3.1 Data, model, and evaluation
A typical pretraining dataset is a mixture of different sources (e.g., GitHub, arXiv, CommonCrawl, etc.) with distinct sampling weights assigned to each constituent. We simplify this reality by considering a fixed document collection, which is a customized version of the DCLM dataset (li2024datacomplm). The original 4T token DCLM-baseline split contains roughly 80% duplicates, as reported by zyphradedup. Therefore, we begin with the de-duplicated dataset, which consists of 769B tokens. We clean the raw Zyphra de-duplicated data by normalizing repeated line breaks, removing long URL links, and fixing malformed Unicode characters. For efficiency reasons, we cap the context window of the synthesizer-tuning (3) step at 8,192 tokens. As a result, we additionally filter out the documents whose length is above 4,096 tokens, allowing both $d_{1}$} and $d_{2}$} to fit into the context window in the worst case when both documents are 4,096 tokens long. After all the de-duplication, cleaning, and filtering procedures, we end up with a collection of 582M high-quality documents ${\mathcal{D}}_{\text{pretrain}}$} totaling 482B tokens. We use the notation $|{{\mathcal{D}}_{\text{pretrain}}}|$ to denote the number of documents in the pretraining dataset and $\|{{\mathcal{D}}_{\text{pretrain}}}\|$ to denote the total number of tokens.
@@
We use the Llama 3 transformer architecture (llama3) to model the probability $p_{\theta}$ with the notable exception of implementing a QK-norm on top of the existing design, which we empirically find to stabilize training. Our resulting model is a 3B-parameter 26-layer transformer model with a hidden dimension of 3,072. Each layer employs grouped query attention with 24 query heads and 8 key/value heads. To validate the scalability of SBP, we also train a 6B-parameter model with 32 layers, a hidden dimension of 4,096, 32 query heads, and a feedforward dimension of 13,056 (detailed in Table 12). The position embedding is RoPE (rope) for queries and keys, with frequency 5e+5. The feedforward network (FFN) has hidden dimension 8,064, and we apply prenorm to both the attention and FFN blocks. For tokenization, we implement a customized BPE tokenization with a vocabulary size of 49,152. To match the 8,192 context window design for synthesizer-tuning we have mentioned, we use context window 4,096 for pretraining, so that every document in ${\mathcal{D}}_{\text{pretrain}}$} can fit into the context window.
@@
To assess the pretraining capability of LM, we measure pretraining test loss and general world knowledge benchmarks. We evaluate held-out test perplexity (exponential of negative log-probability) on 1) OpenWebText2 from EleutherAI (gpt2); 2) Narrative understanding with LAMBADA (paperno2016lambada) and 3) Broad domain multiple-choice with MMLU (hendrycks2020measuring). We evaluate QA accuracy on 4) Hard scientific reasoning with ARC‑Challenge (clark2018think); 5) Easy scientific reasoning with ARC‑Easy (clark2018think); 6) Scientific QA with SciQ (Welbl2017CrowdsourcingMC); 7) Common sense reasoning with Winogrande (sakaguchi2021winogrande); 8) Reading comprehension with TriviaQA (joshi2017triviaqa); 9) Openbook QA with WebQS (berant-etal-2013-semantic). We directly evaluate the pretrained model with either zero-shot or few-shot prompts. Although MMLU is more commonly known as a QA benchmark, we find that evaluating MMLU accuracy for weak models yields a highly non-smooth readout. As a result, for each MMLU test question, we prepend the question with a 5-shot example of QA pairs and postpend it with the correct answer. Then, we treat each such sample as a text corpus and evaluate LM’s perplexity on such a text sample. Empirically, we find that this perplexity-based MMLU correlates well with MMLU accuracy when the underlying model is large enough to yield a stable readout, and also delivers smooth performance changes for smaller models. Note that those benchmarks are known to improve significantly with instruction finetuning (flann). However, we stick to our data-constrained setup and do not introduce any additional data that may confound the comparison.
@@ 3.2 Compute-matched comparsion
We propose a compute-matched experimentation framework to rigorously compare SBP against two natural references: a repetition baseline where we repeat ${\mathcal{D}}_{\text{pretrain}}$} multiple times to utilize the available training compute and an oracle upper bound that enables the model to access as many unique documents as possible. Operationally, we control the training compute by controlling the total tokens seen during training, which is proportional to the training FLOPs given a fixed batch size and context window. We validate SBP across three different settings:
@@
  • 1T-scale (3B): We also consider a larger scale closer to frontier model training, where we cap the training compute at 1T tokens and data access at $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B tokens.
  • 1T-scale (6B): To validate SBP on larger models, we additionally train a 6B-parameter model with the same 1T token budget and 50B unique data access.
@@
Since the compute budget typically exceeds the total number of unique tokens $\|{{\mathcal{D}}_{\text{pretrain}}}\|$, a natural baseline to use the additional compute is to repeat ${\mathcal{D}}_{\text{pretrain}}$} over multiple epochs. By design, in both 200B-scale and 1T-scale, we repeat the pretraining dataset ${\mathcal{D}}_{\text{pretrain}}$} 20 times to exploit the available compute budget. In practice, when the pretraining dataset comes from a mixture of different sources, higher-quality documents can be seen as many as 30 times during pretraining, while lower-quality texts may appear only once. muennighoff2023scaling systematically evaluates the repetition baseline as a proposal to scale LMs under data constraints and finds that repeating ${\mathcal{D}}_{\text{pretrain}}$} up to 4 times yields nearly no performance degradation compared with having access to unlimited fresh data, but after around 40 times, repetition yields rapidly diminishing returns. Therefore, our choice of 20 times repetition with compute-matched comparison strikes a reasonable balance between efficient experimental execution and exhausting all possible performance gains from a fixed ${\mathcal{D}}_{\text{pretrain}}$} via repetition.
@@
Besides showing improvement against the repetition baseline, we also evaluate an oracle upper bound with unlimited data access. The motivation behind this is to contextualize the numerical improvement delivered by SBP. As we shall see in the next section, because different benchmarks respond differently to data size changes, SBP can deliver an improvement as large as 3.74% on some benchmarks but only 0.14% on others (Table 1). Also, as performance on LM benchmarks tend to scale logarithmically (owen2024predictablelanguagemodelbenchmark; kaplan2020scalinglawsneurallanguage) against data improvement, the numerical difference quickly caps out as we move from the 200B scale to the 1T-scale. By introducing this oracle upper bound, we can contrast the SBP improvement against this “oracle” improvement.
@@
For both the repetition baseline and oracle upper bound at all scales, we use a batch size of 2,048 and a context window of 4,096, resulting in a throughput of 8M tokens per step. We apply a cosine learning rate scale with a 5% warmup to a peak learning rate of 1e-2, followed by subsequent decay to 5e-5 towards the end. Under this setup, pretraining costs 11K v5p-TPU hours at 200B-scale, 59K v5p-TPU hours at 1T-scale (3B), and 265K v5p-TPU hours at 1T-scale (6B). For a clean comparison, we adhere to this hyperparameter throughout the paper, including the SBP experiment presented next.